|
EUROPEAN SOUTHERN OBSERVATORY
La Silla - Science Operation Department
|
Summary of IR activities in 2004
|
Version: 1.0 2005/Jan/22 Ivo Saviane |
V. Ivanov needs more training with TIMMI2: give him survival guide?
- Personnel: Malvina Billeres finished her duties in LSO
On Oct 1st, Olivier Hainaut is leaving to Paranal (after having been
appointed deputy to the Paranal head of Science Operation)
- Change of f/35 electronics
- Status:
Requirement specs for the modification of the M2 Chopper Servo and
the Rotator Motor repair/modification issued by U. Weilenman (Feb
29)
Conceptual review done on April 21, now final design being prepared
A review of the f/35 upgrade project is scheduled for July 9, meanwhile
J. Valenzuela produced the design descritption 1.0 (LSO-DSD-ESO-63106-0001)
FDR of the f/35 upgrade project will be most likely on Aug. 17
- Migration to IRACE with PC-Linux
- Quad Timing
- Status:
The 'quad' readout mode was tested, and it failed. The image quality
was vary bad, and the limiting flux > 200 MJy in N1 (< 40 with the
single readout mode). More tuning of IRACE needs to be done.
- lunar occultation mode
- Status:
o Lunar occultation mode tested successufully by U. Kaufl. From Ulrich
himself: "Dear Colleagues please have a look at the enclosed
link; Bringfried has done a nice movie; so as you can see, lunacr
occultation with TIMMI2 works; thanks to all of you who helped"
Movie is at http://www.tls-tautenburg.de/research/stecklum/T2_LO.mpg
But: xxx
- TIMMI2 user manual archivable/Latex version
- TIMMI2 pipeline document
- Problem with the px scale in FITS headers (reported by Malvina). Bernarnd
Lopez is working on it, Malvina will check and report
- PC00010001 etc. keywords should be replaced by PC1_1 etc. keywords.
An AP has been opened.
- o From John: <<At the start of Period 74, the SofI
sotpl CMM module from which the Instrument Package is created was
"hacked" so that it more or less behaves like a
standard TSF module under APR2004, by which I mean... 'make clean
all man install install_ip' works completely through to installation
of the IP on te DHS machine and moreover compiles the IP against the
ohsls module so that execution time calculations are included.>>
- UWE made the guideline for calibration of the polarizer
- VLTSW APR2004
- Wrote a template (plesaux) that: takes roughly 20 points around the
zero position, where supposedly the fixed polarizer (N+POLA) and the
rotating polarizer are crossed and give minimum counts at DITs of
6 ms in N-0.2" with sky simulator in. This will be run at
each setup, and at regular intervals during the TIMMI2 time, so we
can see if the polarizer is loose or not.
- IP 1.24 published (sent to S. Marteau)
- Since a user had complained in the EoM report, that there are no OBs
for the faint TIMMI2 astrometric standards, to be used for accurate
pointing during the night, such OBs were made by IS and archived under
CMM (version: 0.3 date: Thu Jun 17 06:59:28 MESZ 2004)
- Templates:
- list of polarization angles set as tsf parameter;
- jitter returning at original position;
- fast exit in spectro acquisition template;
- the nodding position is properly indicated in the OS panel;
- Dichroic has been cleaned
- A technical intervention (Valenzuela+Weilenman) was carried out on
June 4-11, the following were done:
- The polarizer could have moved in the barrel, the clamping was not
very well and it was now glued with one small dot into the barrel.
- The rotating polarizer gives minimum flux at 78.64 deg w/respect to
the fixed polarizer. This _must not_ change from now on. If it does,
then the polarizer is loose again.
- The objective lens for imaging in L and M has a damaged AR coating,
and it was discovered that this has been since at least 2 years (H.U.
Kaufl, priv. comm.). UWE sent a picture, and it looks really awful:
practically the whole surface of the objective is affected.
- The array mask was out of place, probably exposing the multiplexer
to light. This could be the culprit of the lower sensitivity! At the
moment the three objectives (N-0.2", N-0,3" and
Q-0.3") are aligned such that approx. 5 rows at the upper
edge of the detector are vignetted, in order to prevent the MUX from
being illuminated. This should be kept through each and every run!
- The entrance window was cleaned. The flatfield now looks better, but
the window did not look really that bad.
- Wavelength calibration was checked (both N- and Q- bands), it's within
0.1 pix. of what it was since last calibration in May 2003.
- ICS status display has been checked, and necessary changes will be
requested from Juan Carlos Guzman.
- The PREMA temp. sensor Ch#7 is definitely bad. <<We
do not have a spare, and this instrument is not being produced anymore,
so we have maybe to live with this. Fortunately it's not an important
data point...>>
- Window Cleaning. Cleaning has been assigned to Alain. Note that the
window disassemblin and re-assembling is a critical operation, one
must follow the procedure from hukaufl, see Appendix A.
- Temp. Sensor Repair. (Baffle)
- Detector: Attach Spring Washers on lead block.
- Alignment of the Polarizer. Alignment of the polarizer, uweilenm should
be there, so he can show how to do it.
- dichroic re-coating. see Appendix B for details
- DPR keywords fixed, now data can be visible in the ESO archive
HIERARCH ESO DPR DID |
= |
ESO-VLT-DIC.TIMMI_OS
HIERARCH ESO DPR CATG |
So HIERARCH ESO DPR DID and HIERARCH ESO DPR TYPE were added, and
ACQ for HIERARCH ESO DPR CATG were changed to OTHER
- all polarimetry templates were changed. Now the accept a list of wiregrid
polarizer position angles, while previously only 4 fixed angles were
acquired.
- The [SIV] filter was finally tested. A line was added in the modetab
list with the appropriate parameters. Observations of a planetary
nebula with known [SIV] flux and of Jupiter were carried out,
and we conclude that the filter is ready for scientific use.
- The 'reverse voltage' in IRACE was tuned in order to increase the
sensitivity. See voltages.dat. V=-7.5 gave the best limiting flux
(39.7 mJy), so it has been set to this value now.
- SWC (jguzman) installed the latest version of TIMMI2 software, which
yields a new status panel. To open the new panel, in the User Menu
select "ICS Status Panel" or from a console type:
tiipanStatus. The panels shows temperatures from prema and Lakeshore,
the presure, and different instrument status signals (this last is
not cabled yet). All parameters can be displayed as numeric fields
or graphs vs. time
- After fixing the position of the detector mask, the sensitivity reached
excellent values. In October, a sensitivity of *17mJy/10s/hr*
was reached in N1, which confirms that the instrument is competitive
with VISIR.
- Report on the polarizer by Gaspare and Paul: /sci/facilities/lasilla/sciops/doc/timmiPol.html
- Armin Silber <asilber@eso.org> sent the mechanical drawing of the
polarizer. After inspection, G. Ihle points out that the possible
problems could be with the motor clutches (outside T2), or the motor-pol.wheel
clutch inside T2 (drawing part near #39).
- Occultation template is not working: It needs proper user requirements
to Paul before going ahead. At the moment only takes one burst frame
(object).
- The "_seq" templates for polarimetry cycle A-B-A-B
rather than A-B-B-A like the others. They work fine but might be slightly
less time efficient.
- Problem with exposure failure (IRACE I/O). had to restart DCS software.
It is happening ~ 1/night (3 nights statistics). 5-10
minutes downtime at each occurrence.
- IP installed at the telescope is 124.04, IP distributed by Garching
to V.A.s is 124.06. Execution times as measured at the telescope are
wrong, but they are calculated right by the users who use 124.06.
It might be that 124.06 implements the changes to the EVM I gave to
Stephane Marteau while 124.04 doesn't.
- Sensitivity pipeline ported from w3p6drl to w3p6off where now runs
almost fine. GLO could not get the sensitivity plot. May be an IDL
problem related with the porting of the pipeline on w3p6off. However
the table is correctly updated with the new values.
- PC-Linux - Sparc: PC is still unstable. Day time tests performed.
- Polarizer motor and filter wheel failed few times during the run.
Under test by Javier.
- The cooling of the detector is still not ideal. It was thought that
tightening and spring loading of the bolts on the detector cold finger
could improve cooling, however during the intervention it was found
that the screws were pretty tight already. So now the lowest achieveable
temperature on the detector is 6.5 K.
- lots of tests of the polarization mode were done during two slots
of tech nights. A report was sent earlier, the short story, with upgrades,
is:
The flux in N1 vs. pol. angle is a cosinusoidal of period 177.3 deg
and zero point at -7.687 deg. The N1 band is the worse case, with
inst. polarization of ca. 30%. According to Ulrich, this is due to
the inclination of the filter. Generally, the inst. polarization increases
for longer wavelengths, with the exception of light at N1, which is
the most polarized. This is basically confirmed by N-band spectroscopy.
U. Kaufl thinks that the inst. pol. is created at the filter wheel
(however that would not explain the results from spectropolarimetry).
With respect to the less than 180deg period, Ulrich talked to Pat
Roche, who said that "it is a feature of wire-grid analyzers,
that they are nont optically uniform and rotating the device produces
not only a modulation of intensity due to the polarisation effect,
but there is also simply a modulation of the optical efficiency."
According to Ueli, the grism rotation is not controlled by an encoder:
"Only thing we do, is counting steps from the init position.
Of course we pretend not to lose steps. But this can easily be checked
with respect to the fixed polarizer (we are always assuming, that
nothing slips or jumps)." So a bad calibration of the encoder
cannot explain the weird period introduced by the polarizer.
- complete the characterization of the Js band filter
- complete he reduction of the spectroscopic sensitivity
- Fix tool that selects telluric standards (see above)
- Find way to compute statistics (efficiency) automatically
- Make web page to keep track of projects
- Document new warmup/cooldwon procedure
- Make web page to keep track of projects
- Find way to compute statistics (efficiency) automatically
- Several tests of the polarizer have been suggested by Ueli, after
the problems reported in November. To be done asap.
- Put online the UWE guideline for calibration of the polarizer (and
all other procedures, by the way)
- Convert wtimmi into Linux machine
- Put all transmission curves on the web. Note that seems like there
are no t. curves for L and M filters, but L and M filters in TIMMI2
are basically the same as in ISAAC (hukaufl).
- The anti-reflection coating of the objective for L and M imaging is
in very bad conditions, we might think of replacing it. Indeed H.U.
Kaufl states that it's a relatively straightforward matter, he'll
check with A. Morwood when back to Garching after the VISIR run
- The connectors for the f/35 top unit (going from the chopper to the
rack in the 3.6m cage) have no protections (or have loose protections
that fall). So the pins can be bent or broken: it happened this last
run, making us lose a full tech night. It is then urgent to make safe
covers for all the plugs.
- From ukaufl, <<monitor the polarizer, by measuring
the relative angles of the "scientific" polarizer
and the "technical" polarizer on the filter wheel.
It can not be excluded, that the "scientific" polarizer
is loose in its mount. The check with the "technical"
polarizer is a means to check. In case of problems one can later remeasure
the position angle of the technical polarizer with the Brewster-window
set-up in the lab. In that sense position angle on the sky can be
re-calibrated, if necessary.>>
- update update start-up check list accordingly to have the new temperature/pressure/signal
status panel always opened in one of the available workspaces.
- Make checklist for tech/setup nights